OpenStudy (anonymous):

I need help with this! I've been trying to solve it and I keep getting confused. Here's the problem: s=lw+wh+hl solve for h. I need the steps on how to do it also

OpenStudy (anonymous):

+- s = w + wh + h does this help ?

OpenStudy (nubeer):

s-Iw = wh+hI s-Iw = h(w+I) h = (s - Iw) / (w+ I)

OpenStudy (anonymous):

s = lw + wh + hl Subtracting lw from both sides s - lw = lw + wh + hl - lw wh + hl = s - lw Taking h as common term. h ( w + l) = s - lw \[h = \frac{ s - lw }{ w + l }\]

OpenStudy (anonymous):

mm i looked at it if it was an absolute value | |

OpenStudy (anonymous):

I think I get it now. Thanks guys

OpenStudy (anonymous):

You are most welcomed.